OpenStudy (anonymous):

0, 4, 8 find an equation with these solutions

OpenStudy (saifoo.khan):

\[x^3-12 x^2+32 x \]

OpenStudy (saifoo.khan):

\[(x-0)(x-4)(x-8)\]Multiply them.

OpenStudy (asnaseer):

generally an equation of the form: \[(x-c_1)(x-c_2)...(x-c_n)=0\]where all the "c" terms are constants means that the "c" terms are all the solutions of tis equation. using this you should be able to work out the answer you want.

OpenStudy (anonymous):

saifoo.khan your answer is wrong

OpenStudy (saifoo.khan):

No, it can't be.

OpenStudy (saifoo.khan):

chk your question

OpenStudy (anonymous):

loll i did and it says wrong

OpenStudy (asnaseer):

look at what I typed out - that should give you a BIG hint

OpenStudy (asnaseer):

sigh...it should be: (x-0)(x-4)(x-8)=0 which simplifies to: x(x-4)(x-8)=0

OpenStudy (saifoo.khan):

And then finally to answer i wrote.

OpenStudy (anonymous):

ur right saifoo... my homework much just be stupid

OpenStudy (asnaseer):

which expands to: \[x^{3}-12x^{2}+32x=0\] you omitted the "= 0" on the RHS.

OpenStudy (saifoo.khan):

Online hw sucks at formatting. >_<

OpenStudy (anonymous):

oh duh me! i forgot to set it equal to zero